If the perimeter of a rectangular house is 25 1/3 yards, and the length is 22 feet. What is the width?
what are the steps involved, im not just looking for the answer, I just want to learn this. Thanks Kevin.

Answer by edjones(8007)

25 1/3 = 76'
p=2(L+w)
2(22+w)=76
44+2w=76
2w=32
w=16'
.
Ed